OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-5, -8) and is perpendicular to 10x – 6y = -11

OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-5, -8) and is perpendicular to 10x – 6y = -11

OpenStudy (precal):

Find the slope of 10x-6y=-11 Take the opposite reciprocal of the slope, that will make the new equation perpendicular to it.

OpenStudy (anonymous):

y=-3/5 x+b ===> -8=15+b==>b=-23===<y=-3/5 x -23

OpenStudy (radar):

As precal stated find the slope of the line represented by 10x-6y=-11 rearranging to the slope intercept form: -6y=-10x-11 dividing by -6 to get: y=(10/6)x+11/6, m the slope is 10/6 or5/3 Take the negative reciprocal of the slope to get the slope of a line that is perpendicular which is -(3/5). Now use the point slope formula to get the equation of the perpendicular line: -3/5 = (-8-y)/(-5-x) 15+3x=-40-5y -5y=3x-55 y=(3/-5)x-11 or y=-(3/5)x-11